### v4

# Neuromorphic detection and cooling of microparticles in arrays

Neuromorphic detection and cooling of microparticles in arrays
Micro-objects levitated in a vacuum are an exciting platform for precision sensing due to their low dissipation motion and the potential for control at the quantum level. Arrays of such sensors would offer increased sensitivity, directionality, and in the quantum regime the potential to exploit correlation and entanglement. We use neuromorphic detection via a single event based camera to record the motion of an array of levitated microspheres. We present a scalable method for arbitrary multiparticle tracking and control by implementing real-time feedback to simultaneously cool the motion of three uncoupled microscale objects.
Introduction
Modern technology relies on mechanical sensors, from accelerometers in mobile devices [1] to wearable health monitors [2]. As sensors are miniaturized, their surface-to-volume ratio increases and they dissipate more energy via their thermal contacts and through surface strain [3], limiting their performance. By levitating nano- or micro-particles under ultra-high vacuum conditions, using optical, electrical or magnetic fields [4, 5], one creates a mechanical oscillator with remarkably low dissipation [6, 7]. Force sensitivities of yoctonewtons [8, 9] and torque sensitivities at the level [10] have been achieved, with levitating sensors achieving quality factors in excess of [6], motivating researchers to use these systems to search for dark matter [11, 12, 7] and gravitational waves [13, 14].
The control of levitated particles allows the exploration of a wide range of fundamental science [4, 5], and the demonstration of cooling to the ground state of an optical potential [15, 16, 17, 18, 19, 20] opens the door to macroscopic quantum physics [21, 22, 23, 24]. An emerging frontier in this field is the study of arrays of particles, which in the quantum regime would allow generation of entanglement [25] and tests of quantum gravity [26, 27]. Interactions have been observed between pairs of levitated nanoparticles in optical [28, 91, 30, 87], electrodynamic (Paul) [32] and magnetic [33] traps. Detecting and controlling multiple particles in vacuum has so far involved either single particle control with sympathetic cooling [34, 87, 32] or small arrays of optical traps [35, 36]. Some applications will require the control of arrays of tens, or even thousands, of levitated particles [11].
We use neuromorphic imaging for the control of arbitrary particle arrays across a wide field-of-view. Neuromorphic sensors are highly efficient detectors which mimic neurobiological information gathering [37, 38, 39, 40, 41]. Dynamic vision sensors (DVS) are neuromorphic sensors which mimic the retinal response [42], detecting changes across a threshold on each pixel in an array asynchronously to produce a stream of events [43] ideally suited for object tracking [44]. Together with event-driven processing algorithms [45, 46, 47], DVS can achieve microsecond temporal resolution, sub-millisecond latency and high dynamic range detection (>) with minimal data output at low power consumption [38, 48, 41]. Therefore, DVS are highly suited to high-speed and real-time applications requiring low-power in environments with uncontrolled light levels such as in robotics [49], autonomous driving [50] and space flight [51], as well as finding applications in microscopy [52, 53] and astronomy [54]. In this work, we use an event based camera (EBC) with integrated DVS to monitor the motion of an array of levitated particles with a bandwidth high enough to demonstrate real-time simultaneous feedback control of multiple uncoupled particles in an array.
We implement cold damping feedback [94] to cool the motion of the levitated particles [56, 95], a technique with demonstrated quantum ground state cooling capabilities [16, 17, 18]. We cool a single microsphere to a temperature of a few Kelvin and single degrees-of-freedom of multiple particles. In this work, the number of objects we can simultaneously control is only limited by the number of output channels in our feedback electronics. This single-device method for cooling and controlling particles in arrays is readily scalable due to the low data output of neuromorphic detection. Arrays of cooled micro-sensors will lead to enhanced signal-to-noise sensing through sensor fusion [58, 59, 60], enable force gradient sensing [25], and provide a larger interaction area without increasing the mass of the sensor [11]. Due to the low power consumption of neuromorphic detectors our presented methods are ideal for integration into chip-scale technology [93].
Results
Neuromorphic imaging of levitated particle arrays
We levitate arrays of charged m diameter silica microspheres in a linear Paul trap under vacuum conditions [62, 63, 64, 65] (see Methods) and record their motion using an event based camera (EBC) [89], see Fig. 1(a). The charged particles form a stable array due to the Coulomb repulsion between them. Our particular Paul trap geometry (Fig. 1(b)), and the particles’ distribution of charge, means that our particle arrays are non-uniform.
The EBC uses a neuromorphic DVS, which is an array of independent pixels featuring contrast detectors which output an event [38] in response to light levels on the pixel crossing a user-defined threshold. Pixels which do not experience the required level-change output no signal, removing the data-redundancy present in conventional cameras [38], while allowing the full sensor to be used at all times, sometimes referred to as a dynamic region-of-interest. The EBC hardware bundles asynchronous events into equal-length frames, and uses filters to identify objects within its field-of-view [44], after which a proprietary generic tracking algorithm (GTA) tracks the motion of each object independently. We have previously demonstrated single-particle object tracking with an SNR above 35 dB, and for a more detailed analysis of EBC performance in the context of levitated microparticles see [89].
The EBC allows the tracking of multiple objects with a high bandwidth and a linear scaling in data output with the number of tracked objects, as compared to a rapid increase in data volume with increased region-of-interest in a conventional camera. In our system, with fixed magnification, tracking a single particle at 1 kHz using the entire field-of-view (mm2) of the EBC uses kB s-1 as compared to 64,800 kB s-1 using a standard CMOS camera (Thorlabs CS165MU/M) at the same frame rate and field-of-view. This means that the EBC can track many hundreds of particles before the data volume becomes comparable to standard camera technology. By not having to restrict the region-of-interest, the EBC can track objects dispersed over several hundred micrometres whilst retaining high spatial resolution [89]. The low data volume leads to a correspondingly low power consumption, see Supplemental Materials S5 for more details.
In Fig. 1(c) we show an example of a single neuromorphic sensor being used to detect and track 10 levitated particles simultaneously. The EBC identifies each object and tracks it in 2D (illustrated by the coloured boxes), assigning each one a stable identification number allowing us to process the position data of each particle independently. The linear Paul trap defines the coordinate system for the levitated particles, Fig. 1(b). The image on the EBC has a coordinate system , Fig. 1(a). The axes are projected at onto the axis, and the and axes are parallel, see Supplementary Materials S1. This projection allows us to detect all three axes of motion of the levitated particles.
In Fig. 2(a) we reconstruct the motion and relative position of four levitated particles obtained from the 1 kHz tracking algorithm of the EBC, which can be accessed in real-time. Particles levitated by a Paul trap undergo harmonic motion. In Fig. 2(b) we generate the power spectral density (PSD) from the motion of each particle in the array of four to analyze their motion in frequency space. Each particle has a different charge-to-mass ratio (see Supplementary Materials S3) and is levitated in a different part of the confining field, meaning each particle has different frequency modes of oscillation, which are well separated under vacuum conditions. Below, we use this fact to independently cool the motion of multiple particles.
The four particles are aligned along the axis. Motion of the charged particles in this direction leads to coupling between them via the Coulomb interaction [33, 67, 87, 32]. This is evident via the collective modes seen in Fig. 2(b). By controlling the separation between the particles, we can control the coupling strength, see Supplementary Materials S2 for further details. For the multi-particle cooling presented below, we work in a regime where the coupling between the modes is too small to measure, and perform cooling along the direction.
The signal-to-noise for the different particles varies due to non-uniform illumination and varying coupling to electronic noise. It ranges from 3-30, with exact values given in Supplementary Materials S3. All data in this work is taken at gas pressures of 2.0-mbar unless otherwise stated.
Single particle cold damping using neuromorphic imaging
There are many reasons why it is desirable to control the energy of a levitated particle. Although the sensitivity of a levitated sensor does not increase through cooling [68], rapid damping of the motion increases the stability and measurement bandwidth of the system. Reduction of the particle energy to the ground-state of the levitating potential [15, 17, 69] opens up a toolbox of quantum control [5] and sensitivity enhancement [70, 71] mechanisms.
Cold damping is a feedback method whereby a force proportional to the velocity of an oscillator opposes its motion. Depending on the phase of the feedback force relative to the motion, this method damps (cools) or amplifies (heats) the oscillator. When cooling, input and output noise of the feedback electronics limits the ultimate temperature.
We process the position data from the EBC using a field programmable gate array (FPGA) to generate a feedback signal proportional to velocity, with variable gain and phase. This signal is filtered and applied to a control electrode, see Fig. 1(a).
Figure 3(a) shows the PSD of a particle’s measured motion along the axis as it is cooled via cold damping. The shape of the PSD from the in-loop detector is given by [94, 95]:
where , is the particle’s mass, is the mode frequency, is the bath temperature and is the total momentum damping rate on the particle’s motion, with being the calculable momentum damping rate due to the pressure of the surrounding gas [4] and being the additional damping controlled by the feedback gain, which is feedback-phase dependent, with the uncontrollable phase delay caused by electronics and data processing. The term is the feedback circuit noise that is modeled as white noise. When is large the feedback can introduce extra noise and leads to heating, as discussed in [95, 72, 93]. The parameters , and can be extracted from a measured PSD by fitting equation (Single particle cold damping using neuromorphic imaging) to the data. Due to voltage noise from the amplifiers driving our Paul trap, the equilibrium temperature of our particles without cooling ranges from , depending on their charge and spatial location in the trap, hence we express temperatures as a ratio. The equilibrium temperatures for all particles in this paper are given in Supplementary Material S6.
According to the equipartition theorem, the temperature of a levitated particle experiencing cold damping is [95, 87, 73]:
Experimentally, the temperature of each mode can be extracted from equation (2) after fitting the measured PSD over the corresponding resonance peak with equation (Single particle cold damping using neuromorphic imaging) [95]. In Fig. 3(b) we show the effect of increasing on the temperature of a single mode of a single particle.
The cooling depends on the phase between the feedback signal, which is proportional to velocity, and the detected motion of the particle. Equation (2) is fit to experimental data as is varied in Fig. 3(c), with and as free parameters. The fitted value of is , noting that one period of phase delay does not significantly effect cooling for an underdamped oscillator [92], see Supplementary Materials S4. The value of obtained by fitting the data in Fig. 3(c) with equation (2) agrees with the value obtained at the same feedback gain by fitting the data with equation (Single particle cold damping using neuromorphic imaging), (Hz, Hz respectively.
Finally, in Fig. 3(d) we show the variation in temperature with by reducing the gas pressure, with and fixed at around Hz and , respectively. At a pressure of mbar, we reach the noise floor of our system, indicated by the grey region, at a temperature corresponding to , representing 17 dB of cooling. The optimal can be derived from equation (2), see Supplementary Materials S7. To further improve cooling one can improve particle illumination and imaging, decrease the noise in the levitation electronics, and replace the GTA of the EBC with an optimized tracking algorithm. Object tracking has the potential to track levitated microparticles at the shot-noise limit [75].
Simultaneous cooling of microparticles in an array
Our neuromorphic imaging system tracks the motion of every object it identifies. We are able to process this information, and make a feedback loop for each degree-of-freedom that is detected. Each feedback loop consists of a dedicated FPGA and set of analogue filters, and we are limited in our experiment to three loops in total. We stress that this is not a limitation of detection or processing power, simply the number of FPGA outputs available to us. For each degree-of-freedom the phase and gain of each feedback loop must be optimized, and filters must be set accordingly.
In Fig. 4(a) we cool two orthogonal degrees-of-freedom (the and oscillation modes) of a single particle. We are sensitive to all three degrees-of-freedom due to the angle our imaging system makes to the principal axes of the trap, see Supplementary Methods S1. The geometry of our Paul trap allows the control of all degrees-of-freedom with a single electrode.
We optimize and fix the feedback parameters, then lower the background pressure to reduce , hence lowering the temperature . The temperature is limited not by our noise floor, but by imperfect filtering pumping energy from the feedback signal into the mode, since it is close in frequency to the mode. At low pressures this causes the particle to become unstable, preventing us from further lowering the pressure. Cross-talk between particle modes would also limit the ultimate cooling temperature, which could be resolved by better design of the feedback electrodes [76].
In Fig. 4(b) we extend our cooling to the mode of two separate particles. As we lower the pressure, the temperature of each mode drops, reaching dB and dB of cooling. Since only high-pass filters are used when cooling in the direction (see Methods), unfiltered noise from one particle is able to heat the uncooled modes of the other. At lower pressures this causes particle instability and prevents further cooling.
In Fig. 4(c) we cool the mode of three different particles, with the corresponding PSDs shown in Fig. 4(d). The issue of imperfect filtering is more pronounced when dealing with more particles, as there are more modes of the system overlapped with the unfiltered noise. We still achieve better than 7 dB of cooling. Filtering can be improved either by separating the particle modes and applying band-pass filtering, or through the use of phase locked loops [77, 78]. For the data in Figs. 4(b-d) we adjust the particle spacing via the Paul trap voltages until the Coulomb interaction is weak enough such that there are no coupled modes, see Supplementary Materials S2. Collective modes in particle arrays can be cooled via sympathetic cooling in the limit that the feedback damping rate is smaller than the coupling strength between the modes [87, 32].
The EBC used in this study has a sensor size of pixels, with each particle image occupying pixels (the coloured boxes in Fig. 1(c)) and having a motional amplitude of 4 pixels. As long as the centre of the particles are separated by approximately 60 pixels, the particles can be individually tracked. Hence, without changing our imaging system we could simultaneously track of order 500 particles with this EBC. Considering the fact that object tracking allows for sub-pixel resolution [79, 52], by changing the magnification of the imaging system this EBC would be capable of simultaneously tracking at least 2000 levitated microparticles, with a correspondingly high data volume.
Discussion
We have presented a scalable method for the detection and control of microparticles levitated in an array using a single neuromorphic detector. Neuromorphic imaging is ideally suited to this task, due to its natural affinity with detecting the motion of multiple objects [44, 80] and low data-transfer rate [38]. The tracking speed in our work is limited by the proprietary tracking algorithm of the EBC. Commercial neuromorphic imaging sensors, such as the one used in this study, transfer data from the sensor to the camera hardware at GHz rates [48]. The development of custom algorithms has enabled object tracking at 30 kHz by working with the asynchronous data streamed from a DVS using only 4MB of RAM on a standard 2.9 GHz Dual Core CPU [52]. For a fixed frame rate the data volume is fixed regardless of the particle motion frequency. If the frame-rate is increased data is transferred more rapidly from the camera and the data volume and bandwidth increases. By pushing above 100 kHz, neuromorphic sensors would be suitable for feedback cooling optically levitated particles to the quantum ground state of motion [17], considering the shot-noise limited potential of object tracking [79, 75]. This would require custom tracking algorithms and interfacing the sensor directly with FPGA or neuromorphic processing electronics [38, 81], which would also enable the read-out and control of object alignment and rotation [82, 80].
We believe that the particle control method presented in this work could be extended to an array of order 100 microparticles, (see Supplementary Materials S8). Multichannel FPGA systems with high-quality digital filters are a common tool in research labs. Paul traps are stable at low pressures [83], where the motional frequencies of levitated particles have sub-Hz linewidths [84]. The naturally varying charge-to-mass ratio of charged microparticles, along with application of electric field gradients, will enable the spectral separation of motional modes, making possible single-particle control and cooling even for large arrays. The neuromorphic detection and cooling presented here is independent of the levitation method, as long as there is optical illumination, meaning it is suitable for small dielectrics in optical traps, charged absorptive materials in Paul traps (such as organic material) [95], and magnetically levitated objects.
Since the motion of levitated sensors is well understood, simple machine-learning could be used to optimize all of the feedback parameters [95] in an array, and optimal tracking algorithms used which have been shown to enable quantum-level control [17]. When combined with the low power-consumption of neuromorphic imaging technology (less than 30 mW per tracked particle, see Supplementary Materials S5), and great progress in chip-scale particle levitation [93], integrated devices containing arrays of quantum sensors are closer to being a reality.
Methods
Experimental setup
Figure 5 shows the experimental setup surrounding our linear Paul trap. The trap consists of four parallel cylindrical trapping electrodes forming a square, with two coaxial cylindrical endcap electrodes, one of which we call the control electrode which is used for feedback control. The distance from trap centre to the surface of the -diameter trapping electrodes is , and one opposing-pair are driven with 360 V amplitude at 1 kHz. Of the other pair of the four electrodes, the lower one has a constant voltage of 3V applied to minimize single-particle micromotion. The two endcaps are diameter and separated by , with a slight misalignment along the axis which causes particles to be trapped along a diagonal in the plane. The proximity of the electrodes to the centre of the trap means that a voltage applied to either one will create a field with a significant component in all three axes, enabling 3D control with a single electrode.
Laser induced acoustic desorption (LIAD) [85] is used to launch microparticles into the trap at a pressure of mbar and then the pressure is lowered to carry out the experiments presented in this manuscript. We typically trap particles of positive charge, ranging from to .
To image the particles, 18 mW of laser light of wavelength nm is weakly focused onto the array. The scattered light is imaged onto an event-based camera (Prophesee EVK1 -Gen3.1 VGA (camera sensor: Prophesee PPS3MVCD, pixels)) using a long working distance microscope. The camera is precisely calibrated using a method outlined in detail in [89]. When dealing with multiple particles, calibration is performed via displacing a translation stage on which our imaging system is mounted by a known amount.
Data processing
The data pipeline is shown in Fig. 6. The operation of the EBC is described in detail in [89]. The generic tracking algorithm (GTA) of the EBC outputs 2D position data for each object it detects. The EBC is communicated with via a Python script, which separates this 2D information into two 1D data streams for each object. The script passes each data stream to one of three FPGA systems (Red Pitaya STEMlab 125-14) to output the position of each particle. The FPGA clock is synchronized to the clock of the EBC to ensuring timing consistency. Code on the FPGA computes the velocity from the position data, and adds a variable gain and phase to the signal to generate the feedback signal. There is a latency of 10 ms in this pipeline, see Supplementary Materials S4. EBCs are available with FPGA systems on the camera hardware, which will significantly reduce this latency.
Each feedback signal is filtered with an analogue filter to isolate each frequency component of motion: a high-pass filter for the signal (Wavefonix 3320 HPF 24 dB per Octave) and a low-pass filter for the signal (Wavefonix 2140 LPF 24 dB per Octave). The filtered feedback signals are combined with a summing amplifier, and sent to the control electrode.
Data Availability
The data supporting this article is openly available from the King’s College London research data repository, KORDS, at https://doi.org/10.18742/28069136
Code Availability
The code used in the present work is available from the corresponding authors upon request.
Acknowledgments
We would like to thank Prof David Moore for useful discussions, Dr Ruvi Lecamwasam for assistance with writing FPGA code, and Dr John Dale for bringing neuromorphic imaging to our attention. This work is supported by STFC Grant ST/Y004914/1, EPSRC Grant EP/S004777/1 and ERC Starting Grant 803277.
Author Contributions
YR - experimental design and build, data taking, data analysis. BS - experimental design. RY - data taking. QW - supporting simulations. JP - data analysis. MR - experimental design. JM - experimental design, data analysis.
Competing Interests
The authors declare no relevant competing interests.
Supplementary Materials
In this Supplemental Material we provide technical details on the experiment and data analysis, including: S1 Coordinate systems; S2 Mode identification for multiple particles; S3 Extended information for the four-particle dataset; S4 Latency in the feedback loop; S5 Power consumption of the event based camera; S6 Bath temperatures for all particles; S7 Cooling limit and the noise squashing; S8 Scalability of multi-mode cooling.
Appendix 1 Coordinate systems
Our setup has two coordinate systems: the Paul trap coordinates , and the camera coordinates , as shown in Supplementary Figure 1 and Fig. 1 of the manuscript. The Paul trap coordinates define the oscillation axes of the levitated particles. The -axis is defined along the diagonal between the two trapping electrodes held at a DC voltage (see Methods in the main manuscript), the -axis along the diagonal between the two trapping electrodes with an AC voltage, and the -axis along the axis parallel to the endcap electrodes. In the camera frame, the trap - and -axes are projected onto the camera’s -axis, and the camera’s -axis is parallel to the -axis. Therefore, we can capture the 3D motion of the levitated particles with the camera’s 2D image.
Appendix 2 Mode identification for multiple particles
Here we explain the method of identifying modes when we levitate multiple particles in an array, using a two-particle case study. We trap two charged microparticles separated by m, with the separation controlled by Paul trap voltages. Via neuromorphic detection, we calculate the power spectral densities (PSDs) of the two particles, as shown in Supplementary Figure 2(a). We identify six modes in total. To distinguish these modes, we resonantly drive the Paul trap using a sinusoidal voltage applied to the control endcap electrode, at each of the six frequencies sequentially, and observe the response of the particles using a CMOS camera, as shown in Supplementary Figure 2(b). For the upper particle the centre-of-mass motion frequencies are with Hz, Hz, and Hz and for the lower particle the centre-of-mass motion frequencies are Hz, Hz, and Hz.
To identify collective modes due to interactions between the charged particles, we calculate the cross-spectral density (CSD), which picks-out only the correlated spectral components [86]. As shown in Supplementary Figure 3(a), none of the six modes exhibit significant coupling, since the separation between the particles (m) is too large.
When the particles are brought into close proximity, we do observe interactions between them. We trap another two particles and reduce the separation to approximately m. We again observe six modes: 33.7 Hz and 58.1 Hz along the -axis, 85.2 Hz and 150.9 Hz along the -axis, and 89.4 Hz and 98.2 Hz along the -axis. A CSD analysis in this case is given in Supplementary Figure 3(b), from which we can observe that both -axis modes and both -axis modes are correlated, whilst the -axis modes appear to be uncoupled. This analysis is extended to the four particles in the manuscript below.
Appendix 3 Extended information for the four-particle dataset
Here we investigate the collective modes of the four levitated microparticles case shown in Fig. 2 of the manuscript. As discussed in the Supplementary Section S2, the collective modes can be found via a CSD analysis. We show CSDs for all pairs of particles in Supplementary Figure 4, and conclude that the four levitated particles are coupled along the Paul trap -axis (14 Hz, 25 Hz, 32 Hz and 35 Hz) due to their layout in the trap [cf. Supplementary Figure 1]. It is expected that four coupled oscillators have four coupled modes, but the nature of them is intricate and cannot be simply attributed as common or breathing modes unlike the two particle case [87, 88].
In contrast, there is only evidence of very small correlations along the - and -axes, and only between nearest-neighbour particles. We do not see these modes in Fig. 2 of the manuscript, and if they are real they are below our detection noise-floor. Therefore, we conclude that the motion of the four levitated microparticles is only coupled along the -axis.
We report the signal-to-noise ratios (SNRs) of our detection for each of the four particles in Supplementary Table 1. We can see the strongest SNR is reported for particle P4, and it gradually reduces for particles further away from this. This is because we use a single laser for illumination focusing on P4. In single-particle experiments, we achieve a SNR up to dB [89]. The SNR can be improved in future experiments using a higher-power illumination system which can be expanded to uniformly illuminate many particles. Alternatively, separate beams could be used for each particle, via e.g. a spatial light modulator [90, 91]. The charge to mass ratios of the four levitated particles are provided in Supplementary Table 2, from which we can observe that the four charges are different, since the masses are the same to within 10% (manufacturer specified).
| Particle 1 | Particle 2 | Particle 3 | Particle 4 | |
| SNR (dB) | ||||
| SNR (dB) | ||||
| SNR (dB) | ||||
| SNR (dB) | ||||
| SNR (dB) | ||||
| SNR (dB) |
| Particle 1 | Particle 2 | Particle 3 | Particle 4 | |
| q/m (C/kg) |
Appendix 4 Latency in the feedback loop
We measure the latency of the entire data pipeline by exposing a levitated particle to a voltage impulse and then monitoring the response in the data streamed from the last step in the pipeline (after the FPGA). In Supplementary Figure 5, we present the result averaged over 40 realisations. We find that the response is within 10 ms.
Our data-pipeline is sub-optimal, since data is transferred from the event based camera via a PC to the FPGA. In the future we aim to use an FPGA directly connected to the camera hardware. In such a case, the primary limitation would be the sensor latency (s) and the communication bandwidth between the camera hardware and a multi-channel DAC. However, in our current system:
Latency from the changes in light intensity at each pixel to event output: typical 200 s from the detector manual.
Data Transfer to Computer: The camera uses USB 3.0 at a maximum data-rate of 4 GBit/s – this will never be the limiting case because it’s much larger than the data transfer at maximum event-rate. USB 3.0 has a latency of about 30 s.
The EBC software then tracks the objects – latency in this step is very hard to evaluate, since the process is proprietary.
Python code processes the tracking data for it to be sent to the FPGA, this is simple and can be considered negligible.
The data is transferred via Ethernet and a network switch to our collection of FPGAs, which has a latency of about 300 s.
Data Processing on FPGA: Each FPGA computes velocity, applies gain and a phase shift, and outputs a feedback signal. Since our FPGA runs at 124 Msps, it is reasonable to say the latency is negligible compared to the data transfer delays.
The delay of applied control signal has two effects (following Ref. [92]). Firstly, it deterministically shifts the phase between the mechanical system and the feedback signal, leading to oscillatory behaviour in the correlations and thus a transition between heating and cooling. Secondly, it causes stochastic dephasing of the mechanical motion relative to the feedback signal as the delay becomes large, which reduces the strength of the correlations.
From the experimental results of Ref. [92], correlation functions of the microparticle centre-of-mass motion as a function of delay remain high for short delays (on the order of a few oscillation periods) when the dynamics are underdamped (as in our system). Since our levitated oscillators have oscillation frequencies below 100 Hz, the latency of 10 ms corresponds to a single period of delay, and has little consequence for our cooling protocol.
Appendix 5 Power consumption of the event based camera
From the neuromorphic sensor manual (PPS3MVCD) the static power consumption of the EBC is 26 mW, plus a dynamic power consumption based on sensor activity. In our experiments the event rate is about 500 kevt/s, leading to a power consumption of about 27.6 mW. This is very low compared to a CMOS camera (Thorlabs CS165MU/M, 1.17 W Max @ 34.8 fps Full Sensor ROI) or a high-speed camera (iX Cameras i-SPEED 230, 17 W at 2500 fps).
Appendix 6 Bath temperatures for all particles
Due to voltage noise from the amplifiers driving our Paul trap, the equilibrium temperature of our particles without cooling ranges from approximately K. This temperature varies depending on their charge and spatial location within the trap. Here we model the electrical noise as a white noise bath and assume the particles are trapped in a quadratic potential. Under the assumption that the equipartition theorem holds, the different bath temperatures used in the main text figures are listed in Supplementary Table 3. In Fig. 4(c), P2 is located closest to the trap center and therefore experiences the least electric field noise, resulting in the lowest bath temperature. In contrast, P1 and P3 are farther from the trap center, and their bath temperatures are approximately an order of magnitude higher than that of P2. A similar trend can be observed in Fig. 4(b), where P1, being farther from the trap center, also exhibits a higher temperature than P2.
| Fig3(b) | Fig3(c) | Fig3(d) | Fig4(a) | Fig4(b) | Fig4(c) | |
P1:
P2: |
P1:
P2: P3: |
Appendix 7 Cooling limit and the noise squashing
At a certain pressure (which determines the gas damping rate ), to achieve maximal cooling the optimal feedback rate can be obtained from equation (2) in the main manuscript by taking [93], which gives
with the optimal feedback rate
Note that is -independent.
The relation between the feedback rate and the cooling temperature is shown in Supplementary Figure 6(a), where we plot equation (2) in the main manuscript at three different pressures. The dots in the plot are the optimal feedback rates and the optimal temperatures corresponding to the respective pressure, computed from Supplementary Equations (1) and (2). As we can see, the optimal feedback rate gets smaller as the pressure goes down.
At a certain pressure, a larger feedback rate results in a lower system temperature (hence a better cooling), until it pass the the optimal value in which case the feedback signal instead heats up the system. This is because the increasing amplitude of the feedback signal also amplifies the noise fed back to the particle, preventing further cooling. The corresponding PSD also appears inverted below the noise floor, a phenomenon known as the noise squashing [94]. In Supplementary Figure 6(b) we demonstrate this by setting the feedback rate Hz and reducing the pressure. Correspondingly, the PSDs of the measured motion changes from unsquashed (blue line, ), to mixed (orange line, ), and finally to squashed (green line, ). Furthermore, the squashed shape is asymmetrical due to the delay in the signal [95]. This makes a small change to the fitting PSD function, as shown in the following:
Suppose we consider a time delay in the feedback signal , the equation of motion reads
Take the Fourier transform with and , we get the particle’s detected motion in the frequency domain
We can calculate the PSD in this situation by taking and assume that two noises and are uncorrelated, which gives the PSD of the measured position in the in-loop detector:
This signal delay can cause a small asymmetry that is visible in the squashed PSD from in-loop detection. In this case, the experimental PSD of particle’s measured motion can be correctly fitted by Supplementary Equations (3), as shown in Supplementary Figure 6(b).
Appendix 8 Scalability of multi-mode cooling
Here we discuss the fundamental limit on the number of modes we can cool with the technique presented in our manuscript. We assume is bounded by the bandwidth of our detection Hz and the minimum resolvable linewidth , such that . Then, to estimate the number of modes, we need to find the minimum linewidth .
The contribution to this limit from the experimental setup comes mainly from filtering. Considering the time duration for signal demodulation, there is a bandwidth limit for filters to pick out certain frequency signal which will set a limit to . A PLL has a very narrow filter bandwidth, for example a Zurich Instruments HF2LI has a minimum bandwidth of Hz and can filter 6 peaks at once. Therefore, with PLL it is possible to pick up feedback signal at mHz resolution.
On the other hand, since we use cold damping to cool the particles, the particles inevitably experience an additional damping rate, which broadens the linewidth and makes the modes harder to distinguish. According to Eq. (2) in the manuscript, by solving , we can obtain an optimal at a given pressure (See Supplementary Equations (2)). This optimal would be the dominant factor in determining the scaling number, whose value reduces as pressure goes down.
If we take the above argument as a rule of thumb, we can estimate the number of modes as follows. Suppose we start cooling at mbar in experiment, at which the particle’s linewidth is about Hz, the optimal feedback rate is Hz. This limits the capacity of our setup to be modes, which can be improved at high vacuum.